(1)/(3)y=(1)/(3)x+(5)/(3)
x-3y=5

Answer by Alan3354(69443)

(1)/(3)y=(1)/(3)x+(5)/(3)
x-3y=5
-------------------------------
(1)/(3)y=(1)/(3)x+(5)/(3)
Is it (1/3)y or 1/(3y) ?
------
Making 5/3 (5)/(3) doesn't clarify anything.